OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)? x - y = -1 x - y = -2 x - y = - 5 x - y = - 6 @Mertsj

OpenStudy (rajee_sam):

can you find the slope first? Do you know how to do that?

OpenStudy (anonymous):

I don't know how to with this type of problem.

OpenStudy (dan815):

have your learnt about equation of line y=mx +b where m = slope and b = yintercept

OpenStudy (anonymous):

Yes, but how would that be set up on this problem?

OpenStudy (rajee_sam):

OK they give you two points. The first thing you can do with the given points is find the slope of the line. The points given to you are say \[(x _{1}, y _{1}) \space and \space (x _{2}, y _{2})\] The slope formula is: \[\frac{ (y _{2} - y _{1})}{ (x _{2} - x _{1}) }\]

OpenStudy (rajee_sam):

Go ahead find the slope and tell me.

OpenStudy (anonymous):

Okay. give me a moment.

OpenStudy (rajee_sam):

@unheard are you there?

OpenStudy (dan815):

slope = rise/ run rise over run means, how much is the height changing with respect to the width

OpenStudy (anonymous):

I got 4/4

OpenStudy (dan815):



OpenStudy (rajee_sam):

Too many cooks. I do not want to confuse you. Find me if you still have problem

OpenStudy (dan815):

yep that is right so 4/4 = 1

OpenStudy (rajee_sam):

yes. 4/4 is = 1

OpenStudy (anonymous):

But 1 isn't in any of the choices. Just -1.

OpenStudy (dan815):

rajee sam, let me teach her, i want to teach her geometrically!

OpenStudy (rajee_sam):

Now the slope intercept form of a line is given my "y = mx + b" Where "m" is the slope and "b" is the y- intercept

OpenStudy (dan815):

this child will start writing equations of everything on her own after today

OpenStudy (anonymous):

You guys are confusing me.

OpenStudy (dan815):

lol

OpenStudy (rajee_sam):

OK Dan take care of it I will leave

OpenStudy (dan815):

okay we will call u if shes still stuck!

OpenStudy (dan815):

thanks

OpenStudy (dan815):

okay unheard, are you ready

OpenStudy (anonymous):

Yes.

OpenStudy (dan815):

first tell me, have you seen this equation before y=x

OpenStudy (anonymous):

no.

OpenStudy (dan815):